A liquid with specific gravity of 1.2 fills a volume. If the mass in the volume is 146kg
, what is the magnitude of the volume?
1
Expert's answer
2020-12-14T12:17:08-0500

Solution:

"SG=1.2"

"m=146kg"

"p_{H_2O}=1000( \\tfrac{kg}{m^3}\t)"


"The\; specific\; gravity\; is\; given \; by:"


"SG= \\dfrac{p}{p_{H_2O}}\t= \\dfrac{ \\tfrac{m}{V}\t}{p_{H_2O}}"


"where\; p=\tfrac{m}{V}\; -is \;the\; density\; of\; the\; liquid,\; and\; p_{H_2O}-\;is \;the \;density\; of \;the \;water."


"V= \\dfrac{m}{SG*p_{H_2O}}\t= \\dfrac{146}{1.2*1000( \\tfrac{kg}{m^3}\t)}=0.121\\;m^3"


"Answer:\: 0.121\;m^3"